### 11 Dec 2023 v5

# Cellular Wireless Networks in the Upper Mid-Band

Abstract
The upper mid-band – roughly from 7 to 24 — has attracted considerable recent interest for new cellular services. This frequency range has vastly more spectrum than the highly congested bands below while offering more favorable propagation and coverage than the millimeter wave (mmWave) frequencies. The upper mid-band can thus provide a powerful and complementary frequency range to balance coverage and capacity. Realizing the full potential of these bands, however, will require fundamental changes to the design of cellular systems. Most importantly, spectrum will likely need to be shared with incumbents including communication satellites, military RADAR, and radio astronomy. Also, the upper mid-band is simply a vast frequency range. Due to this wide bandwidth, combined with the directional nature of transmission and intermittent occupancy of incumbents, cellular systems will need to be agile to sense and intelligently use large spatial and frequency degrees of freedom. This paper attempts to provide an initial assessment of the feasibility and potential gains of wideband cellular systems operating in the upper mid-band. The study includes: (1) a system study to assess potential gains of multi-band systems in a representative dense urban environment and illustrate the value of wide band system with dynamic frequency selectivity; (2) an evaluation of potential cross interference between satellites and terrestrial cellular services and interference nulling to reduce that interference; and (3) design and evaluation of a compact multi-band antenna array structure. Leveraging these preliminary results, we identify potential future research directions to realize next-generation systems in these frequencies.
Index Terms:
Upper mid-band, 6G, cellular wireless systems, FR3, satelliteI Introduction
Cellular wireless systems up to the 4G had largely operated in a range of microwave frequencies below . Given the severe spectral shortage in these bands, 5G systems [dahlman20205g, 3GPP38300] introduced new capabilities in the millimeter wave (mmWave) frequencies above [RanRapE:14, AkdenizCapacity:14, shafi20175g]. In the US, individual carriers had spectrum allocations as large as – several times larger than previous allocations in the sub-6 bands. These wide bandwidths, combined with spatial multiplexing capabilities, enabled massive multi-Gbps peak rates. However, extensive measurements have now shown that practical performance is often intermittent [narayanan2021variegated, wei20225gperf] with limited penetration inside [narayanan2022comparative, rochman2022comparison]. This poor coverage is due to the inherently limited range of mmWave signals and their high susceptibility to blockage [Rappaport2014-mmwbook, slezak2018empirical, slezak2018understanding, maccartney2017rapid, zekri2020analysis].
Against this backdrop, the upper mid-band spectrum, roughly from 7-24 , provides a good balance of coverage and spectrum, overcoming the spectral shortage of the sub-6 bands while having favorable propagation and penetration relative to the mmWave bands. For these reasons, the band has recently attracted considerable attention for commercial cellular communications and has been cited by industry as a leading candidate spectrum for next generation (NextG, 5G and beyond) wireless networks [smee2022ten, samsung20226gspectrum, 5Gamerica20223gpproadmap]. The Federal Communications Commission (FCC) Technical Advisory Committee (TAC) has recently identified the upper mid-band as vital for meeting the growing data rate demands [fcc2023preliminary, fcc20236Gworkinggroup]. The 3rd Generation Partnership Project (3GPP), the organization that sets cellular standards, has also begun study of these bands [3gpp38820] – see Section II for more details.
In 3GPP, the upper mid-band is referred to as frequency range 3 (FR3). FR1 initially referred to the traditional spectrum below , which was later expanded to [sathya2020standardization]. FR2 refers to the mmWave spectrum that was introduced in the fifth generation (5G) standard).
I-A Challenges
Development of cellular services in the full range of these bands faces significant challenges:
Need for spectrum sharing and sensing with incumbents: Most importantly, the bands are already used by several vital incumbent services. As we review in Section II-C, commercial satellite services already widely use these bands and have been interested in increased bandwidth allocations, particularly for broadband rural access. The upper mid-band also contains scientifically important spectral lines that are fundamental for radio astronomy as well as vital military RADAR.
Large spectral range: The upper mid-band is simply a vast range of spectrum – well beyond the span of most commercial cellular front-ends that typically operate over a small percentage of the carrier frequencies. This large range, combined with the need to dynamically share spectrum and directionally transmit, demands new front-ends that adaptively and intelligently sense and exploit large bandwidth and spatial degrees of freedom.
Spectral resiliency: Sensing capabilities, combined with the ability to adapt in space and frequency, can also yield improved defenses to external attacks from jammers and signal disruption. Such features can provide robust protection for vital future cellular infrastructure.
I-B Contributions
The broad goal of this paper is to assess the feasibility and potential of agile wideband cellular systems operating in these bands. We will also identify key challenges and areas for research to realize next-generation systems in these bands. Towards this end, our contributions are:
Channel modeling and estimation of multi-band coverage: We briefly review channel measurements in the upper mid-band range and discuss open areas of research (Section III-A). Also, based on current propagation models, we conduct a ray-tracing simulation in a representative dense urban area (New York City) to assess the coverage and capacity of a potential cellular system in the upper mid-band (Section III-B). Our results demonstrate that mobiles that can adaptively select one of multiple bands across the upper mid-band can improve coverage and data rates over mobiles restricted to a single band. This finding argues for multi-band, adaptive systems to gain the full benefits of the frequency range.
Interference with incumbents: We perform an additional ray tracing simulation (Section LABEL:sec:sat_int) to assess the potential interference of terrestrial cellular services on the satellite uplink. The analysis shows that interference from both user equipment (UEs) and base stations (BSs) can be significant, but directional nulling can be effective in mitigating the interference with some loss in the terrestrial network capacity. We also discuss some potential interference issues with passive radio astronomy in the upper mid-band.
Wideband antenna systems: Finally, we present (Section LABEL:sec:antenna) a compact, multi-band antenna array to demonstrate the feasibility of front-ends that operate over the entire upper mid-band in a small form factor.
II Background and Standardization Landscape
II-A Prior developments in the sub-6 and mmWave bands
The interest in the upper mid-band has to be seen in the context of developments of commercial cellular systems in both the sub-7 and mmWave frequency ranges over the last decade. Early mmWave experiments and capacity analyses such as [akoum2012coverage, rappaportmillimeter, RanRapE:14, AkdenizCapacity:14] suggested the possibility of massive data rates in micro-cellular deployments owing to the wide available bandwidths and spatial multiplexing gains available in the mmWave frequencies. Based partially on these and other results, the mmWave bands have emerged as an integral component of the 5G New Radio (NR) specification [dahlman20205g, 3GPP38300]. Commercial mmWave deployments appeared shortly after the release of the speceification, particularly in the US [shafi20175g].
Coverage at reasonable cell densities in these deployments, however, has been an enormous challenge. Several recent measurements in commercial mmWave networks now clearly demonstrate that outdoor coverage can be highly intermittent [narayanan2021variegated, wei20225gperf] with limited penetration indoors [narayanan2022comparative, rochman2022comparison]. MmWave signals are simply blocked by many common building materials such as concrete [Rappaport2014-mmwbook, zekri2020analysis] as well as the human body and other obstacles [slezak2018empirical, slezak2018understanding, maccartney2017rapid]. See also Section III-D below.
Parallel to the developments in the mmWave, significant spectrum was also released in portions of the mid-band frequencies above , including the Citizens Broadband Radio Service (CBRS) band (3.55 to 3.7 ) [fcccbrs] and the C-band [FCC20-22] (3.7 to 4.2 ). These bands were instrumental for both so-called private 5G networks [aijaz2020private] as well as spectrum expansion in many wide area public networks. More recently, there has also been significant discussion of unlicensed bands from approximately 6 to 7 , extending the mid-band further [sathya2020standardization]. In particular, both wireless LAN and cellular services have been considering enhancements to operate in these bands [IEEE80211ax, naik2021coexistence].
Measurements from commercial network monitoring companies, now offer an opportunity to compare the practical performance of these networks. On the one hand, Ookla reported in 2022 [ookla2023mmwaverates] that the tested mmWave networks offer an incredible median downlink throughput, up to seven times higher than systems in C-band. Similarly, in 2021, Open Signal demonstrated [opensignal2021mmwave] that Verizon’s mmWave downlink capacity was at least three times as high as networks relying mostly on mid-band spectrum. Nevertheless, of mid-band networks enabled downlink speeds in the hundreds of Mbps. Moreover, the coverage was much more uniform. For example, Open Signal’s study [opensignal2021mmwave] showed that users connect to mmWave less than 1% of the time.
II-B Recent standardization efforts in the upper mid-band
The relative success of commercial cellular systems in the mid-band, combined with the high data rates, but intermittent coverage, of mmWave has set the stage for interest in the upper mid-band. The hope is to provide the high data rates close to those available in the mmWave range, but with much more uniform coverage. As mentioned in the Introduction, the bands have been identified by industry [smee2022ten, samsung20226gspectrum] and 3GPP has recently formally started study of services in the 7 to 24 band [3gpp38820, 5Gamerica20223gpproadmap], where they are called frequency range 3 or FR3.
Note that, in 3GPP terminology, frequency range 1, or FR1, has been extended from sub-6 GHz to sub-7 GHz to include the frequencies in the 6 to 7 range [3GPP38921]. Nevertheless, many discussions use the term upper mid-band to refer to 6 to 24 .
In 2022, the FCC also began discussion on two bands from 12.2-12.7 [FCC20-443], and 12.7-13.2 [FCC22-352] for possible cellular use. While the first band has now been rejected for cellular services, there is wide recognition of the need for opening the upper mid-band for cellular services. For example, the FCC Technical Advisory Committee (TAC) has recently published a comprehensive survey of spectrum from 7.125 to 24 [fcc2023preliminary] as well as 6G working paper [fcc20236Gworkinggroup]. These analyses emphasize the need for considering larger portions of the upper mid-band for terrestrial cellular services in order to meet the growing data demand.
II-C Incumbency and the need for spectrum sharing
The FCC TAC analyses [fcc2023preliminary, fcc20236Gworkinggroup] also emphasize that a key issue in allocating the upper mid-band is incumbents, particularly commercial satellite services that also need bandwidth. In fact, the rejection of the 12.2-12.7 band proposed in [FCC20-443] for cellular services was largely due to the interference onto ground satellite units. We will perform some simple cellular-satellite interference calculations in Section LABEL:sec:sat_int.
The FCC analysis [fcc2023preliminary] has thus considered several potential models for spectrum sharing in the upper mid-band. Interestingly, spectrum sharing was critical in CBRS and C-Band allocations [mueck2015spectrum, matinmikko2020spectrum]. Spectrum sharing mechanisms in these bands included spectrum access systems SAS and more general licensed spectrum access (LSA) schemes are now being considered in the upper mid-band as well. More dynamic methods for spectrum sharing, as is being considered by 3GPP [3GPP21917, QcomDSS] could also be used.
In short, the upper mid-band is a vast and valuable frequency range for numerous services. How to allocate and share the spectrum between various users will be one of the fundamental design and policy challenges going forward.
III Coverage and Capacity Gains
III-A Prior channel measurements and studies
Aspects of channel propagation in the upper mid-band have been studied for over two decades. For example, [janssen1992propagation, janssen1996wideband] studied performed indoor radio measurements up to in the early and mid-1990s. Significant research has continued. For example, additional indoor measurements can be found in [de20182] (from 2.4 to 61,) and in corridors from 9 to 11, in [batalha2019indoor]. Attenuation measurements for various building materials at 800 MHz to 18 were conducted in [deng2016indoor]. Indoor and outdoor 5G diffraction measurements at 10, 20 and 26 can be found in [rodriguez2014radio]. The work in [kristem2018outdoor] performed wideband outdoor channel measurements at 3–18 and a more recent comprehensive set of outdoor and indoor measurements at 3.3, 6.5, 15 and 28 was conducted in [miao2023sub]. Satellite propagation has also been extensively studied – see, for example, [pan2001high] on earth satellite measurements at .
A key and consistent finding of these channel measurement studies is that the factors that influence large-scale propagation, such as transmission losses, reflectivity, and diffraction, vary considerably over the upper mid-band. As expected, the lower frequencies in the upper mid-band provide more favorable coverage in most indoor and outdoor scenarios. At the same time, higher frequencies offer better bandwidth since bandwidth allocations generally scale with the carrier frequency.
III-B Multi-frequency outdoor capacity gains
One implication of the variability of propagation across the upper mid-band is that cellular systems would ideally have access to bands across the spectrum with real-time band selection. Wideband mobiles could use the higher frequencies with higher bandwidth when coverage is available, and switch to lower frequencies when the higher frequencies are blocked.
Parameters | Values |
Area (m) |
Inter-site distance between gNBs (m) |
Maximum transmit power of gNBs (dBm) |
(from 3GPP TR 38.141 [3GPP38141], Table 6.3.1-1, Local Area BS) |
Number of sectors for gNBs |
Down-tilted antenna angle of gNBs |
UE Noise figure (dB) |
7 |
Field pattern of antenna element |
Taken from 3GPP TR 37.840 [3GPP37840] |
Frequency-dependent parameters | ||||
Frequency [GHz] |
6 |
12 |
18 |
24 |
Bandwidth [MHz] |
100 |
200 |
300 |
400 |
BS URA dimensions |
UE ULA dimensions |
To estimate the potential gain of such a multi-band, frequency adaptive system in the upper mid-band, we consider a dense urban area, i.e., Herald Square in New York City, as shown in Fig. 2a and downlink scenario (gNBUE). In dense urban scenarios, capacity is mostly limited, while providing satisfactory coverage is challenging at high frequencies due to blockage, a key challenge, for example, that has been experienced in the mmWave bands [slezak2018empirical, slezak2018understanding, jain2018driven]. Within this area, we see systems operating at up to four potential frequencies: 6, 12, 18, and 24.
The simulation parameters of the potential cellular downlink systems at these frequencies are shown in Table I. Note that the bandwidth in each frequency scales with the carrier, as is typical in deployments today. As shown in Fig. 2b, the size of the gNB antenna array also scales so that the aperture is approximately constant. Furthermore, in the simulation area, we manually selected the locations of 18 BSs (gNB) on rooftops, corresponding to an inter-site distance (ISD) of approximately , typical for urban microcellular evaluations. Terrestrial UEs are randomly placed outside buildings, and ray tracing was performed using Wireless Insite [Remcom], which has recently been proven fairly accurate by conducting real-world measurements in the mmWave bands, as reported in [de2023ray, remcommpresnt2017modeling].
We compute the wideband SNR for each UE associated with the serving BS. At each UE location, we select one BS serving with the strongest SNR. Fig. 2c shows a coverage map for a single BS deployment as an example. We see that coverage is greatly reduced as frequency increases. Similarly, the curves in Fig. 3 for frequencies 6, 9, 18 and 24 represent the cumulative distribution function (CDF) of the signal-to-noise ratio (SNR) and rate for a hypothetical UE.
To compute the achieved rate (goodput) with consideration of system overheads, we use a standard realistic model [mogensen2007lte] as the follows
where is the bandwidth, is a system bandwidth loss factor to account for overhead and receiver imperfections, and is a maximum spectral efficiency. As indicated in [mogensen2007lte], we adopt for and for .
As shown in Fig. 3a, UEs at lower frequencies experience better SNRs due to benign propagation conditions and reduced noise power using a smaller bandwidth compared to higher frequencies. We plot the CDFs of the data rate for each frequency in Fig. 3b, where the maximum data rates for , , and are specified as , , and , respectively. As expected, the peak rates for UEs are higher at higher frequencies (18 and 24 ) due to wider bandwidths. However, cell edge UEs (for example, UEs up to the bottom 10% percentile of the CDF) suffer a significantly worse rate in the higher frequency bands (18 and 24 ) compared to UEs in the lower frequencies (6 and 12 ). This property is also expected, as propagation is much less favorable for UEs at the edge of the cell at higher frequencies.
The curves labeled “best choice” in Fig. 3b correspond to the rates for the UEs that select the best BS and frequency. They achieve the same peak rates as UEs in the high-frequency bands. At the same time, we observe that the cell edge rate is improved over any individual frequency. These results suggest that cellular systems that adaptively select across a range of bands through the given spectrum are able to obtain peak rates at higher carrier frequencies while improving cell edge rates.
III-C Penetration Loss
In the previous section, we focused on outdoor UEs. Coverage for indoor from outdoor cell sites is called outdoor to indoor (O2I) penetration. To understand the potential for O2I coverage in the upper mid-band, Fig. 4 plots the 3GPP model [3GPP38901, Table 7.4.3-1] for the loss of the O2I path for several common exterior building materials. For each material, the 3GPP model for pathloss is given by
where is the path loss in dB, is the frequency in GHz, and and are linear constants that depend on the carrier frequency and are given in [3GPP38901, Table 7.4.3-1]. Note that the curve in Fig. 4 does not appear linear since the pathloss is plotted on a logarithmic scale. We see that standard glass and wood are relatively permeable throughout the frequency range ( for both materials), while infrared reflecting glass (IRR) is relatively impermeable across all RF frequencies (). Concrete, however, makes a sharp transition from relatively permeable to inpenetrable, precisely in the upper mid-band.
III-D Multi-frequency Indoor Capacity Gains
The 3D models used in the ray tracing have approximate material classifications for the facades of each building. Using the penetration loss models from Eq. 2 combined with the exterior wall classification, Fig. 5 plots the estimated CDFs of the SNR and data rate for indoor UEs. Fig. 5a shows that the SNR for indoor UEs is considerably reduced at higher frequencies due to high penetration loss, as mentioned in the previous section. In this setting, concrete is a dominant exterior wall material and significantly reduces the signal penetration – see Fig. 4. Fig. 5b shows that, for roughly 65% of the users, the data rates at lower frequencies are higher than at higher frequencies. These results indicate a further potential gain of wideband cellular systems in FR3: the lower frequencies can provide valuable indoor coverage behind materials such as concrete while the higher frequencies can opportunistically offer high capacity for outdoor users and selected indoor users with minimal blockage (e.g., indoor users next to non-IRR glass windows).
III-E Effects of Interference
As the last discussion, we consider a full buffer interference scenario where BSs transmit data to their associated outdoor UEs simultaneously, and each UE receives interference signals from other BSs. Note that this is the worst-case scenario to show the lower bound of the system capacity gains. As before, considering ISD, BSs are manually selected and concurrently transmit data to UEs.
Fig. 6 shows the CDF of the signal-to-interference-noise ratio (SINR) and the corresponding data rate calculated using Eq. 1. As shown in Fig. 6a, SINRs are degraded at lower frequencies (6 and ) because the small number of antenna arrays produces wider beamwidths. In contrast, due to directional transmissions with narrow beamwidths, UEs at higher frequencies experience less interference. As a result, we observe in Fig. 6b that data rates at higher frequency are almost uniformly better than at lower frequency.
However, some cautions on this result should be taken: the simulation is only for outdoor users. Due to the high penetration loss, indoor users will still likely require lower frequencies. Moreover, our simulation does not have blockers other than buildings. Practical scenarios, particularly for urban users at the street level, will experience blocking by vehicles and people that can reduce coverage. Nevertheless, an interesting finding from this result is that dynamic frequency selection in the upper mid-band can enhance the coverage and capacity gains when interference levels vary considerably. For example, when systems at lower frequencies become interference-limited, users can be moved to higher frequencies, provided that the propagation is not blocked.
IV Interference with Incumbent Services
IV-A Commercial communication satellites
One of the most vital and growing incumbents in the upper mid-band are commercial communication satellites [huang2020recent]. Fig. LABEL:fig:satbands shows the standard satellite bands. We see immediately that the X, Ku, and K bands – all widely used by satellites – fall entirely in the upper mid-band. In fact, recent FCC reports [fcc2023preliminary, fcc20236Gworkinggroup] indicate extensive use of the bands by various commercial satellite services. Moreover, with the rapid growth of satellite Internet services, there is enormous demand for increased bandwidth, particularly in the 7-15 bands. As an example, Fig. LABEL:fig:satbands shows the bands granted to the commercial satellite provider Starlink in the recent FCC grant [fcc-starlink-2022]. As depicted, the bands in the grant are in both the uplink and downlink, as well as gateway-satellite and terminal-satellite links. 3GPP has also begun considering 5G NR services from satellites to mobile devices [3GPP38811, 3GPP38821, lin20215g], including communication in the S, K, and Ka bands, which fall partly within the upper mid-band. More generally, satellite services are growing rapidly [grandview2023satellite] and if terrestrial cellular services are to be deployed in significant fractions of the upper mid-band, co-existence with satellite services will be crucial.